# DiTASK: Multi-Task Fine-Tuning with Diffeomorphic Transformations

### v2

@@ Appendix E Qualitative Comparison
Figure 10 shows semantic segmentation results on the PASCAL MTL [13] dataset, comparing MTLoRA and DiTASK against the ground truth. DiTASK consistently produces sharper and more accurate segmentations. In the first row, it captures the structure and boundaries of bicycles more precisely, whereas MTLoRA oversmooths the outputs, failing to recover fine details. Similarly, in the third row, DiTASK segments smaller objects, such as zebras, with greater detail, avoiding omissions observed in MTLoRA. For complex indoor scenes, such as rows five and six, DiTASK distinguishes multiple objects effectively and maintains segmentation coherence, whereas MTLoRA generates fragmented outputs. These results highlight DiTASK’s ability to adapt pre-trained weights through diffeomorphic transformations, enabling accurate segmentation across diverse object categories.
Appendix F Additional Related Work
Hard Parameter Sharing and Task Dynamics in Multi-Task Learning. Hard parameter sharing is a widely used approach in multi-task learning (MTL), where most layers of a neural network are shared among tasks, while task-specific layers are restricted to the output heads. This technique, introduced in [7] is computationally efficient but presents challenges due to task interference, where conflicting task gradients degrade performance. Approaches like PCGrad [55] mitigate this by enforcing gradient orthogonality, but they do not always address the full extent of task competition for limited shared parameters. Despite these challenges, task synergies can be harnessed through careful parameter modulation, allowing shared features to benefit related tasks. In this context, methods like our DiTASK enhance positive transfer by preserving crucial pre-trained feature structures, enabling both task-agnostic and task-specific adaptations through diffeomorphic transformations.

### v3

@@ Appendix E Qualitative Comparison
Figure 10 shows semantic segmentation results on the PASCAL MTL [13] dataset, comparing MTLoRA and DiTASK against the ground truth. DiTASK consistently produces sharper and more accurate segmentations. In the first row, it captures the structure and boundaries of bicycles more precisely, whereas MTLoRA over-smooths the outputs, failing to recover fine details. Similarly, in the third row, DiTASK segments smaller objects, such as zebras, with greater detail, avoiding omissions observed in MTLoRA. For complex indoor scenes, such as rows five and six, DiTASK distinguishes multiple objects effectively and maintains segmentation coherence, whereas MTLoRA generates fragmented outputs. These results highlight DiTASK’s ability to adapt pre-trained weights through diffeomorphic transformations, enabling accurate segmentation across diverse object categories.
@@
Appendix F Additional Related Work
Hard Parameter Sharing and Task Dynamics in Multi-Task Learning. Hard parameter sharing is a widely used approach in multi-task learning (MTL), where most layers of a neural network are shared among tasks, while task-specific layers are restricted to the output heads. This technique, introduced in [7], is computationally efficient but presents challenges due to task interference, where conflicting task gradients degrade performance. Approaches like PCGrad [57] mitigate this by enforcing gradient orthogonality, but they do not always address the full extent of task competition for limited shared parameters. Despite these challenges, task synergies can be harnessed through careful parameter modulation, allowing shared features to benefit related tasks. In this context, methods like our DiTASK enhance positive transfer by preserving crucial pre-trained feature structures, enabling both task-agnostic and task-specific adaptations through diffeomorphic transformations.